OpenStudy (anonymous):

d=k/n rewrite the formula to determine how man laps Elena must run around the lake to reach her goal. d=distance around lake k=total number of kilometers n=number of laps Elena runs around the lake Help me ):

OpenStudy (anonymous):

omg marry me!! XDD

OpenStudy (anonymous):

lol your so lucky moderators arent here

OpenStudy (anonymous):

haha no xD

OpenStudy (anonymous):

REJECTED

OpenStudy (anonymous):

answer my math question and I will

OpenStudy (anonymous):

what about me :(

OpenStudy (anonymous):

k = d.n

OpenStudy (anonymous):

answer the questioooooooooooooooooooooooooooooooooooooooooooon

OpenStudy (anonymous):

c: please and thank you :3

OpenStudy (anonymous):

lol its her first question xD

OpenStudy (anonymous):

yeah c: I had such a hard time trying to log in with my facebook DX

OpenStudy (jumperman):

k=d/n I think.

OpenStudy (anonymous):

it wont let me put letters only numbers and it never gave me any numbers to use .-.

OpenStudy (jumperman):

Then it's looking for the answer?

OpenStudy (anonymous):

yes sir c:

OpenStudy (jumperman):

Does it say how many kilometers?

OpenStudy (anonymous):

No all it tells me is Elena is getting in shape for the soccer season. Her coach suggests running a certain number of km each week. Elena runs laps arouna a lake near her house to reach this goal

OpenStudy (anonymous):

and then Here's a formula for this scenario: d= k/n k=km d= distance around lake n= number of laps elena runs around the lake Rewrute the formula to determine how many laps elena must run around the lake to reach her goal.

OpenStudy (jumperman):

There might be a problem with the actual question, is it an online school?

OpenStudy (anonymous):

Product of ( no. of laps and distance around lake) gives the total distance covered Therefore k = n.d

OpenStudy (anonymous):

yeah its treca :0

OpenStudy (anonymous):

What do I do I don't want to get it wrong :c

OpenStudy (jumperman):

Well then if they won't let you put in letters, you will need to contact someone about that.

OpenStudy (anonymous):

I mean if I guess it wrong the first time it gives me another try and it will give me helpo it might give me numbers hold on

OpenStudy (jumperman):

Try that, if it shows letters then tell someone about that.

OpenStudy (anonymous):

well before i was trying to put in the answer i got and it wouldnt let me but now its letting me use letter and it told me n=k/d

OpenStudy (anonymous):

maybe the thing was all jacked up at first

OpenStudy (jumperman):

Try a different question.

OpenStudy (anonymous):

will you help me answer it?

OpenStudy (jumperman):

Absolutely

OpenStudy (anonymous):

okay same question but with numbers this time the lake is 3 kilometers Elena wants to run 12 kilometers this week how many laps around the lake?

OpenStudy (jumperman):

4

OpenStudy (anonymous):

Yess!!!! c:

OpenStudy (jumperman):

She wants to run 12 laps, so 12 divided by 3 equals 4.

OpenStudy (anonymous):

You should just stay here and help me with this study assignment?

OpenStudy (jumperman):

Or we can do it in messages, if you want specific help from me.

OpenStudy (anonymous):

okay!

OpenStudy (jumperman):

Get my message?